OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -2, and -1 + 2i How would I solve this?

OpenStudy (campbell_st):

so the polynomial has 4 roots.. the real roots are x = 4 and x = -2 the complex roots are \[x = -1 \pm i\]

OpenStudy (campbell_st):

so to find the complex roots start with \[x = -1 \pm i\] add 1 to both sides \[x + 1 = \pm i\] square both sides of the equation \[(x + 1)^2 = i^2\] remember i^2 = -1 then \[x^2 + 2x + 1 = -1\] add 1 to both sides of the equation and you will have the quadratic factor of the polynomial... so the polynomial is \[P(x) = (x^2 + 2x +1)(x -4)(x+1)\] just distribute for the equation

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

2i, 5, -5?

OpenStudy (campbell_st):

is that another question...?

OpenStudy (anonymous):

No, is that the correct answers @campbell_st

OpenStudy (campbell_st):

well the polynomial is \[P(x) = (x^2 + 2x + 2)(x -4)(x +2)\] which becomes \[P(x) = (x^2+ 2x +2)(x^2 -2x - 8)\] you just need to multiply the 2 quadratic factors together

OpenStudy (anonymous):

The answer I come up with is not one of the options. X^4-10x^2-20x-16

OpenStudy (mertsj):

f(x)=(x-4)(x+2)[x-(-1+2i)][x-(-1-2i)]

OpenStudy (mertsj):

Multiply the four factors and that will be your answer.

OpenStudy (anonymous):

Okay I got x^4+6.5x^2-26x

OpenStudy (anonymous):

Is that correct?

OpenStudy (mertsj):

f(x)=(x^2-2x-8)(x^2+5)=x^4-2x^3-8x^2+5x^2-10x-40=x^4-2x^3-3x^2-10x-40

OpenStudy (mertsj):

That's what I get from multiplying those 4 factors. Is that one of your choices?

OpenStudy (anonymous):

Yes, I must have done it wrong. Thank you.

OpenStudy (anonymous):

Wait, no.

OpenStudy (anonymous):

what the hell

OpenStudy (mertsj):

If 4is a root, then x-4 is a factor. If -2 is a root then x-(-2)=x+2 is a factor If -1+2i is a root then -1 - 2i is also a root If -1+2i is a root then x-(-1+2i) is a factor If -1-2i is a root then x - (-1-2i) is a factor

OpenStudy (mertsj):

So now we have 4 factors of the polynomial: (x-4)(x+2)[x-(-1+2i)][x-(-1-2i)] (x-4)(x+2)=x^2-2x-8 [x-(-1+2i)][x-(-1-2i)]=x^2-x(-1-2i)-x(-1+2i)+(-1+2i)(-1-2i)= x^2+x+2xi+x-2xi+5=x^2+2x+5

OpenStudy (mertsj):

So now we must multiply (x^2-2x+8)(x^2+2x+5)= x^4+2x^3+5x^2-2x^3-4x^2-10x+8x^2+16x+40= x^4+9x^2+6x+40

OpenStudy (mertsj):

Is that one of your choices?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

What are the answer choices?

OpenStudy (anonymous):

a. f(x) = x4 - 7x2 - 26x - 40 b. f(x) = x4 - 3x3 - 8x2 - 13x - 40 c. f(x) = x4 + 6.5x2 - 26x - 40 d. f(x) = x4 - 3x3 + 8x2 + 13x + 40

OpenStudy (mertsj):

I made a typo when I multiplied: (x^2-2x-8)(x^2+2x+5)=x^4+2x^3+5x^2-2x^3-4x^2-10x-8x^2-16x-40= x^4-7x^2-26x-40

OpenStudy (mertsj):

Before I used x^2-2x+8 instead of x^2-2x-8

OpenStudy (anonymous):

Ah I see.

OpenStudy (anonymous):

Thanks.

OpenStudy (mertsj):

yw